### v1

@@ 1 Introduction
In the past decades, the study of topological phases of matter has been a significant subject [1, 2, 3, 4, 5, 6]. Due to their profound significance and extensive potential for next-generation devices with ultralow power dissipation, topological materials have emerged as a frontier in fields of condensed matter physics as well as material science. Recent research has seen a surge in classifying and realizing topological states in crystalline materials, thanks to progress in symmetry-based theoretical frameworks and computational methods [7, 8, 9, 10, 11]. Various catalogues of topological materials, such as nonmagnetic and magnetic topological electronic materials [12, 13, 14, 15, 16, 17, 18], topological phononic materials [19, 20, 21, 22, 23, 24, 25], and topological superconductors [26, 27, 28, 29, 30, 31, 32, 33, 34], have been successively established. More recently, by further utilizing symmetry arguments, researchers have comprehensively constructed the effective models of all magnetic space groups [35, 14, 36, 37] and established the encyclopedia of emergent quasiparticles in three-dimensional crystals [38, 39], strongly facilitating the development of topological physics. However, the current focus mainly lies in investigating equilibrium states of topological physics, while studies on topological physics in non-equilibrium scenarios are still in the infancy. With the increasing achievements of topological states and topological materials, performing investigations of quantum control such as creation and manipulation of topological states is an imperative task.
@@
Light-matter interaction is an important approach to dynamically modulate material properties on ultrafast timescales, enabling the creation of exotic non-equilibrium topological states that are otherwise not possible in the equilibrium cases [40, 41, 42, 43, 44, 45, 46]. Among various mechanisms of light-matter interaction, of particular interest is the concept of Floquet engineering. Within the framework of Floquet theory [47, 48, 49, 50], periodic light field transfers Bloch energy bands of crystalline solids to periodic Floquet-Bloch sidebands through multiphoton absorption or emission. Thus, light driving offers a mean to manipulate electronic structures, enabling a great potential to control electronic topology out of equilibrium in materials. By appropriately selecting incident light that matches the target system, Floquet engineering can provide a wide range of pathways for dynamically manipulating topological states and even inducing topological phase transitions. For instance, light can gap out the Dirac cone in graphene or drive a band inversion in semiconductor quantum wells, thereby developing the notion of a Floquet topological insulator (FTI) [43, 51, 52, 53, 54]. The topologically nontrivial band gap can also be induced from avoided crossings of photon-dressed Floquet sidebands via the optical Stark effect [55, 56, 57, 58].
@@
The Floquet topological phases are strongly dependent on the drive frequency, amplitude, and polarization of incident light, exhibiting high tunability. It has been demonstrated that irradiation of circularly polarized light (CPL) can break the time-reversal ($\mathcal{T}$) symmetry, and thus quantum anomalous Hall (QAH) insulators with a non-zero Chern number can be obtained under irradiation of CPL [41, 59, 42, 60, 61, 62, 63]. Unlike the QAH effect that is present in magnetic materials, the Floquet QAH effect does not require initial magnetism and can generally be present in light-irradiated magnetic and nonmagnetic materials [64, 65]. Besides, Floquet engineering can give rise to Chern flat bands with tunable large Chern numbers in twisted systems irradiated by CPL [66, 67, 68]. Inspired by the development of conventional FTI with the dipole polarization, it is found that periodic driving can further lead to Floquet higher-order topological insulators with multipole polarization [69, 70, 71, 72, 73, 74, 75, 76]. Beyond gapped topological phases, tailoring topological semimetallic phases with first-order or higher-order topology through light irradiation have also been intensely studied [77, 78, 79, 80]. Through changing the propagation or polarization direction of incident light, light control of symmetry breaking can be accordingly achieved, and periodic driving via light irradiation offers a fascinating avenue to realize desired gapless topological fermions. The typical examples are the Floquet Weyl semimetals (WSMs) with highly controllable Weyl nodes, which can be generated in light-irradiated topological insulators [80], Dirac semimetals (DSMs) [81, 82, 83, 84], nodal-line semimetals (NLSMs) [85, 86, 87, 88].
@@
On the other hand, disorder plays an important role on the observation of topological edge or surface transport in a realistic topological system [89, 90, 91, 92, 93, 94, 95], and even can induce a phase transition from a topologically trivial insulator to a topological Anderson insulator (TAI) phase [89]. In periodically driven systems, it has been demonstrated that disorder can induce topological phases that go beyond the well established paradigm of static disorder-induced topological phases [94, 95, 96]. The topological phase arises from the interplay of disorder and periodic driving, dubbed a Floquet TAI. Experimentally, the Floquet topological sates as well as their emerged exotic properties can be captured by time-resolved and angle-resolved photoemission spectroscopy (TrARPES) or time-resolved transport measurements [97, 98, 99, 100, 101]. For instance, the Floquet sidebands and CPL-induced gap of topological surface states were observed in three-dimensional topological insulator Bi2Se3 by using TrARPES [99, 98]. Through the application of ultrafast time-resolved transport measurements using a laser-triggered ultrafast photoconductive switch, the light-induced anomalous Hall effect was confirmed in CPL-driven graphene [102]. Recently, the Floquet band engineering in a semiconductor has achieved important progress in experiments, such as pseudospin-selective Floquet sidebands in black phosphorus [103], and optical control of valley polarization in semiconductors MoS2 [104] and BN [105].

### v2

@@ 1 Introduction
In the past decades, the study of topological phases of matter has been a significant subject [1, 2, 3, 4, 5, 6]. Due to their profound significance and extensive potential for next-generation devices with ultralow power dissipation, topological materials have emerged as a frontier in fields of condensed matter physics as well as material science. Recent research has seen a surge in classifying and realizing topological states in crystalline materials, thanks to progress in symmetry-based theoretical frameworks and computational methods [7, 8, 9, 10, 11, 12, 13]. Various catalogues of topological materials, such as nonmagnetic and magnetic topological electronic materials [14, 15, 16, 17, 18, 19, 20], topological phononic materials [21, 22, 23, 24, 25, 26, 27], and topological superconductors [28, 29, 30, 31, 32, 33, 34, 35, 36], have been successively established. More recently, by further utilizing symmetry arguments, researchers have comprehensively constructed the effective models of all magnetic space groups [37, 16, 38, 39] and established the encyclopedia of emergent quasiparticles in three-dimensional crystals [40, 41], strongly facilitating the development of topological physics. However, the current focus mainly lies in investigating equilibrium states of topological physics, while studies on topological physics in non-equilibrium scenarios are still in the infancy. With the increasing achievements of topological states and topological materials, performing investigations of quantum control such as creation and manipulation of topological states is an imperative task.
@@
Light-matter interaction is an important approach to dynamically modulate material properties on ultrafast timescales, enabling the creation of exotic non-equilibrium topological states that are otherwise not possible in the equilibrium cases [42, 43, 44, 45, 46, 47, 48, 49, 50]. Among various mechanisms of light-matter interaction, of particular interest is the concept of Floquet engineering. Within the framework of Floquet theory [51, 52, 53, 54], periodic light field transfers Bloch energy bands of crystalline solids to periodic Floquet-Bloch sidebands through multiphoton absorption or emission. Thus, light driving offers a mean to manipulate electronic structures, enabling a great potential to control electronic topology out of equilibrium in materials. By appropriately selecting incident light that matches the target system, Floquet engineering can provide a wide range of pathways for dynamically manipulating topological states and even inducing topological phase transitions. For instance, light can gap out the Dirac cone in graphene or drive a band inversion in semiconductor quantum wells, thereby developing the notion of a Floquet topological insulator (FTI) [47, 55, 56, 57, 58, 59]. The topologically nontrivial band gap can also be induced from avoided crossings of photon-dressed Floquet sidebands via the optical Stark effect [60, 61, 62, 63].
@@
The Floquet topological phases are strongly dependent on the drive frequency, amplitude, and polarization of incident light, exhibiting high tunability. It has been demonstrated that irradiation of circularly polarized light (CPL) can break the time-reversal ($\mathcal{T}$) symmetry, and thus quantum anomalous Hall (QAH) insulators with a non-zero Chern number can be obtained under irradiation of CPL [45, 64, 46, 65, 66, 67, 68]. Unlike the QAH effect that is present in magnetic materials, the Floquet QAH effect does not require initial magnetism and can generally be present in light-irradiated magnetic and nonmagnetic materials [69, 70]. Besides, Floquet engineering can give rise to Chern flat bands with tunable large Chern numbers in twisted systems irradiated by CPL [71, 72, 73, 74]. Inspired by the development of conventional FTI with the dipole polarization, it is found that periodic driving can further lead to Floquet higher-order topological insulators with multipole polarization [75, 76, 77, 78, 79, 80, 81, 82, 83, 84]. Beyond gapped topological phases, tailoring topological semimetallic phases with first-order or higher-order topology through light irradiation have also been intensely studied [85, 86, 87, 88]. Through changing the propagation or polarization direction of incident light, light control of symmetry breaking can be accordingly achieved, and periodic driving via light irradiation offers a fascinating avenue to realize desired gapless topological fermions. The typical examples are the Floquet Weyl semimetals (WSMs) with highly controllable Weyl nodes, which can be generated in light-irradiated topological insulators [88], Dirac semimetals (DSMs) [89, 90, 91, 92, 93], nodal-line semimetals (NLSMs) [94, 95, 96, 97].
@@
On the other hand, disorder plays an important role on the observation of topological edge or surface transport in a realistic topological system [98, 99, 100, 101, 102, 103, 104], and even can induce a phase transition from a topologically trivial insulator to a topological Anderson insulator (TAI) phase [98]. In periodically driven systems, it has been demonstrated that disorder can induce topological phases that go beyond the well established paradigm of static disorder-induced topological phases [103, 104, 105, 106]. The topological phase arises from the interplay of disorder and periodic driving, dubbed a Floquet TAI. Experimentally, the Floquet topological sates as well as their emerged exotic properties can be captured by time-resolved and angle-resolved photoemission spectroscopy (TrARPES) or time-resolved transport measurements [107, 108, 109, 110, 111, 112, 113]. For instance, the Floquet sidebands and CPL-induced gap of topological surface states were observed in three-dimensional topological insulator Bi2Se3 by using TrARPES [109, 108]. Through the application of ultrafast time-resolved transport measurements using a laser-triggered ultrafast photoconductive switch, the light-induced anomalous Hall effect was confirmed in CPL-driven graphene [114]. Recently, the Floquet band engineering in a semiconductor has achieved important progress in experiments, such as pseudospin-selective Floquet sidebands in black phosphorus [115], and optical control of valley polarization in semiconductors MoS2 [116] and BN [117].